# Evidence of anisotropic three-dimensional weak-localization in TiSe2 nanoflakes

### v1

@@ Abstract
TiSe2 is a typical transition-metal dichalcogenide known for its charge-density wave order. In this study, we report the observation of an unusual anisotropic negative magnetoresistance in exfoliated TiSe2 nanoflakes at low temperatures. Unlike the negative magnetoresistance reported in most other transition-metal dichalcogenides, our results cannot be explained by either the conventional two-dimensional weak localization effect or the Kondo effect. A comprehensive analysis of the data suggests that the observed anisotropic negative magnetoresistance in TiSe2 flakes is most likely caused by the three-dimensional weak localization effect. Our findings contribute to a deeper understanding of the phase-coherent transport processes in TiSe2.
@@ I Introduction
In this article, we present our study of the magnetotransport properties of exfoliated TiSe2 nanoflakes. The MR data of TiSe2 flakes were measured down to mK in a 3He cryostat. Surprisingly, an unusual anisotropic NMR was observed in both perpendicular and in-plane magnetic fields. It was found that the NMR curves measured under in-plane fields were significantly broader than those measured under perpendicular fields, indicating a strong anisotropy of the NMR. Such an unusual anisotropic NMR deviates substantially from the theoretical expectations of both the 2D WL and Kondo effects, which are commonly associated with NMR at low temperatures. A detailed analysis of the data suggests that the observed anisotropic NMR in TiSe2 nanoflakes presumably originates from the 3D WL effect. Our findings contribute to a better understanding of the phase coherent transport processes in TiSe2.
@@ II.2 Device Fabrication and Transport Measurements
The longitudinal resistivity and Hall resistivity of TiSe2 flakes were measured in a Hall-bar geometry using lock-in amplifiers, with an AC excitation current of 3 A at 17.78 Hz. The measurements were performed in a 3He cryostat with a base temperature of 260 mK.
@@ III.1 Temperature Dependence of Resistivity
The curves of TiSe2 flakes behave similarly to those of bulk TiSe2 crystals [1]. As plotted in Fig. 1(b), upon cooling down from room temperature, first increases until reaches a maximum at K. It then decreases with further lowering of temperature, exhibiting a metallic behavior. The most prominent feature of the curves is the anomalous resistivity peak centered around K, which has long been a focal point in previous studies. Early work speculated that the resistivity peak might be related to the excitons in the CDW ordering [15, 16, 17]. Recently, Watson et al. proposed that the it could be explained by a simple two-band model [18], without invoking the CDW order in TiSe2.
@@ III.2 Magnetotransport
At K, the Hall resistivity [] of TiSe2 flakes remains temperature-independent, as illustrated in Figs. 3(a) and (b), further confirming that the carrier density of TiSe2 remains constant at low temperatures. The slightly curved curves suggest the presence of multiple types of charge carriers in TiSe2, consistent with the results from ARPES measurements [10].
@@ III.3.1 The Kondo Effect
where is the elementary charge, , , , and correspond to the 2D carrier density and mobility of the electron and hole bands, respectively. Here, the electron band represents the 3D-like electron pocket at the point, while the hole band is associated with the 2D-like hole pocket at the point [22]. The fitting results are plotted as dashed lines in Figs. 3(a)-(d).
After subtracting the background, the net resistivity change due to the out-of-plane NMR is obtained, as plotted together with the data in Figs. 3(e) and (f). Clearly, for both devices, the curves are much broader than the curves, indicating a high degree of anisotropy in the NMR. Such anisotropic NMR is inconsistent with the expectations of the Kondo effect, which predicts an isotropic NMR independent of field direction [20, 21]. Therefore, the possibility of the Kondo effect can be safely excluded.
@@ III.3.2 Two-Dimensional Weak Localization
In previous work, the NMR in TiSe2 flakes was attributed to the 2D WL effect [14]. However, in our experiment, the NMR was observed under both perpendicular and in-plane magnetic fields, making the explanation of 2D WL unlikely.
@@ III.3.3 Anisotropic Three-Dimensional Weak Localization
Here, and denote the phase coherence lengths in perpendicular and in-plane directions, respectively, and is the anisotropy parameter. The function is the Hurwitz zeta function, expressed as
@@
To determine whether the observed NMR can be well described by the 3D WL thoery, we analyzed the low-field data from devices A and B using Eqs. (4) and (5). Here, it is important to note that the anisotropy parameter appears in both Eqs. (4) and (5) and therefore should not be taken as an independent parameter when fitting Eqs. (4) and (5) to the and curves measured at the same temperature. However, in this experiment, the and curves were not measured at exactly the same temperature. Therefore, during the fitting, we aimed to collapse the data extracted from the and curves onto a common trend line to ensure consistency.
@@
Fig. 5(f) shows the data extracted from the fitting. As decreases, increases monotonically, whereas shows no clear trend. At low temperatures, reaches about 100 nm, comparable to the thickness of the TiSe2 flakes in both devices, indicating that our system lies at the boundary between 3D and 2D in terms of diffusive transport. The temperature dependence of below 1 K follows the relation , as depicted by the dashed line in Fig. 5(f).
Theoretically, the exponent in the relation is expected to take values of , 1 and 0.75 for systems in which the inelastic scattering is dominated by electron-phonon interaction, EEI in the clean limit and EEI in the dirty limit, respectively [19]. Clearly, the exponent obtained in this experiment does not match any of these theoretically predicted values. Although a similar value of has been reported in the 3D WL of nitrogen-doped ultrananocrystalline diamond films [28], the exact origin of this deviation remains unclear. Further research is needed to fully elucidate the discrepancy.
@@ IV Conclusion
In summary, we observed an unusual anisotropic NMR in exfoliated TiSe2 nanoflakes. Unlike the NMR commonly seen in other 2D materials, our results cannot be explained by either the 2D weak localization effect or the Kondo effect. We tentatively attribute the observed NMR to the anisotropic 3D WL effect. Our work provides valuable insights into the phase-coherent transport processes in TiSe2 and paves the way for further research in this field.

### v2

@@ Abstract
TiSe2 is a typical transition-metal dichalcogenide known for its charge-density wave order. In this study, we report the observation of an unusual anisotropic negative magnetoresistance in exfoliated TiSe2 nanoflakes at low temperatures. Unlike the negative magnetoresistance reported in most other transition-metal dichalcogenides, our results cannot be explained by either the conventional two-dimensional weak localization effect or the Kondo effect. A comprehensive analysis of the data suggests that the observed anisotropic negative magnetoresistance in TiSe2 flakes is most likely caused by the three-dimensional weak localization effect.
@@ I Introduction
In this article, we present our study of the magnetotransport properties of exfoliated TiSe2 nanoflakes. The MR data of TiSe2 flakes were measured down to mK in a 3He cryostat. Surprisingly, an unusual anisotropic NMR was observed in both perpendicular and in-plane magnetic fields. It was found that the NMR curves measured under in-plane fields were significantly broader than those measured under perpendicular fields, indicating a strong anisotropy of the NMR. Such an unusual anisotropic NMR deviates substantially from the theoretical expectations of both the 2D WL and Kondo effects, which are commonly associated with NMR at low temperatures. A detailed analysis of the data suggests that the observed anisotropic NMR in TiSe2 nanoflakes presumably originates from the 3D WL effect.
@@ II.2 Device Fabrication and Transport Measurements
The longitudinal resistivity and Hall resistivity of TiSe2 flakes were measured in a Hall-bar geometry using lock-in amplifiers, with an AC excitation current of 3 A at 17.78 Hz. No obvious overheating caused by excitation current was observed [15]. The measurements were performed in a 3He cryostat with a base temperature of 260 mK.
@@ III.1 Temperature Dependence of Resistivity
The curves of TiSe2 flakes behave similarly to those of bulk TiSe2 crystals [1]. As plotted in Fig. 1(b), upon cooling down from room temperature, first increases until reaches a maximum at K. It then decreases with further lowering of temperature, exhibiting a metallic behavior. The most prominent feature of the curves is the anomalous resistivity peak centered around K, which has long been a focal point in previous studies. Early work speculated that the resistivity peak might be related to the excitons in the CDW ordering [16, 17, 18]. Recently, Watson et al. proposed that it could be explained by a simple two-band model [19], without invoking the CDW order in TiSe2.
@@ III.2 Magnetotransport
At K, the Hall resistivity [] is almost independent of temperature, as illustrated in Figs. 3(a) and (b), further confirming that the carrier density of TiSe2 remains constant at low temperatures. The slightly curved curves suggest the presence of multiple types of charge carriers in TiSe2, consistent with the results from ARPES measurements [10].
@@ III.3.1 The Kondo Effect
where is the elementary charge, , , , and correspond to the 2D carrier density and mobility of the electron and hole bands, respectively. Here, the electron band represents the 3D-like electron pocket at the point, while the hole band is associated with the 2D-like hole pocket at the point [23]. The fitting results are plotted as dashed lines in Figs. 3(a)-(d), and more details are available in the Supplemental Material [15].
After subtracting the background, the net resistivity change due to the out-of-plane NMR is obtained and plotted together with the data in Figs. 3(e) and (f). The procedure for the classical-background subtraction is presented in the Supplemental Material [15]. Clearly, for both devices, the curves are much broader than the curves, indicating a high degree of anisotropy in the NMR. Such anisotropy is inconsistent with the expectations of the Kondo effect, which predicts an isotropic NMR independent of field direction [21, 22]. Therefore, the possibility of the Kondo effect can be safely excluded.
@@ III.3.2 Two-Dimensional Weak Localization
In previous work, the NMR in TiSe2 flakes was attributed to the 2D WL effect [14]. However, in our experiment, the NMR is observed under both perpendicular and in-plane magnetic fields, as shown in Fig. 2. The presence of in-plane NMR strongly points to a 3D origin, rendering the explanation of 2D WL unlikely.
@@
We also note that similar large values of , which far exceeds the expected value for 2D weak antilocalization (WAL), have been reported in several bulk topological semimetals [26, 27, 28]. As in our case, these observations cannot be explained by conventional 2D WAL and instead point to a 3D origin.
@@ III.3.3 Anisotropic Three-Dimensional Weak Localization
Here, and denote the phase coherence lengths in the perpendicular and in-plane directions, is the anisotropy parameter, and the function is expressed as
@@
To determine whether the observed NMR can be well described by the 3D WL thoery, we analyzed the low-field data from devices A and B using Eqs. (4) and (5). Here, it is important to note that the anisotropy parameter appears in both Eqs. (4) and (5) and therefore should not be taken as an independent parameter when fitting to the and curves. However, in this experiment, the and curves were not measured at exactly the same temperature. Therefore, during the fitting, we aimed to collapse the data extracted from the and curves onto a common trend line to ensure consistency.
@@
Fig. 5(f) shows the data extracted from the fitting. As decreases, increases monotonically, whereas shows no clear trend. At low temperatures, falls in the range of 10–20 nm, which is significantly smaller than the flake thickness of approximately 100 nm. This indicates that, in terms of diffusive transport, the TiSe2 flakes should be regarded as 3D systems, consistent with the presence of 3D WL.
The extracted values of are comparable to those typically obtained in other materials, whereas the values of are noticeably smaller. Such a pronounced anisotropy in phase coherence length may originate from both the anisotropic effective mass and scattering rates of electrons in TiSe2, although the exact mechanism remains to be clarified in future studies. If our interpretation is correct, a dimensional crossover from 3D to 2D WL should occur when the flake thickness is reduced below 20 nm. We plan to investigate this transition in future work.
The in-plane phase coherence length reaches about 100 nm at low temperatures and follows a temperature dependence of below 1 K, as shown by the dashed line in Fig. 5(f). Theoretically, the exponent in the relation is expected to take values of , 1 and 0.75 for systems in which the inelastic scattering is dominated by electron-phonon interaction, EEI in the clean limit and EEI in the dirty limit, respectively [20]. Clearly, the exponent obtained in this experiment does not match any of these theoretically predicted values. Although a similar value of has been reported in the 3D WL of nitrogen-doped ultrananocrystalline diamond films [32], the exact origin of this deviation remains unclear. Further research is needed to fully elucidate the discrepancy.
@@ IV Conclusion
In summary, we observed an unusual anisotropic NMR in exfoliated TiSe2 nanoflakes. Unlike the NMR commonly seen in other 2D materials, our results cannot be explained by either the 2D WL effect or the Kondo effect. We tentatively attribute the observed NMR to the anisotropic 3D WL effect. Our results provide direct experimental signatures of phase-coherent transport processes in TiSe2 and may be helpful for advancing the understanding of phase-coherent transport phenomena in other layered TMD materials with similar anisotropic characteristics.
@@ Acknowledgements.
- [15] See Supplemental Material for details regarding the magnetoresistance data under different currents, estimation of uncertainty in resistivity calculation and subtraction of the classical magnetoresistance background.
@@
- [26] Z. Hou, Y. Wang, G. Xu, X. Zhang, E. Liu, W. Wang, Z. Liu, X. Xi, W. Wang, and G. Wu, Transition from semiconducting to metallic-like conducting and weak antilocalization effect in single crystals of LuPtSb, Appl Phys Lett 106, 102102 (2015).
- [27] A. Laha, S. Malick, R. Singha, P. Mandal, P. Rambabu, V. Kanchana, and Z. Hossain, Magnetotransport properties of the correlated topological nodal-line semimetal YbCdGe, Phys. Rev. B 99, 241102 (2019).
- [28] S. Roy, B. Das, A. Ghosh, A. Bera, C. K. Barman, M. Mondal, A. Alam, and P. Mandal, Chiral anomaly and weak antilocalization effects in the Weyl semimetal EuAuSb, Phys. Rev. B 110, 85145 (2024).
@@
Supplemental Material
Appendix S1 Comparison of Magnetoresistance Data under Different Currents
To rule out the possibility of overheating caused by the excitation current, we compared the magnetoresistance data measured at base temperature using three different excitation currents: 1 A, 3 A and 5 A. As shown in Fig. S1(a), the magnetoresistance curves for 1 A and 3 A are closely matched, whereas the curve for 5 A appears noticeably flatter. These results indicate that a 3 A excitation current does not cause overheating in our TiSe2 samples. Additionally, the noise level at 3 A is significantly lower than that at 1 A. Therefore, we select a 3 A excitation current for our main measurements.
Appendix S2 Estimation of Uncertainty in Resistivity Calculation
The length-to-width ratio of the devices was calculated using , where is the distance between the midpoints of the two voltage (potential) contacts, and is the width of the TiSe2 flake, as illustrated in Figs. S1(b) and (c). The selected TiSe2 flakes have well-defined geometries, which minimize the uncertainty in estimating the length-to-width ratio due to sample shape. Although the presence of voltage contacts may perturb the current distribution and introduce some uncertainty in determining the effective length, this uncertainty can be reasonably estimated by taking and as the upper and lower bounds for the length, respectively, as indicated in Figs. S1(b) and (c). For both devices, the maximum relative uncertainty in the length-to-width ratio is approximately , which would correspondingly result in an uncertainty in the calculated resistivity , as well as in the prefactor . However, this level of uncertainty does not affect the validity of our main conclusions.
Appendix S3 Fitting Details of the Classical Magnetoresistance Background
To reliably obatin the classical parabolic background in the curves, we fitted the two-band Drude model simultaneously to the high-field ( T) and data. According to the two-band Drude model, the longitudinal and transverse resistivity are respectively given by
and
where , , and are carrier density and mobility of the two bands. Note that here , are sign-carrying variables, where a positive value indicates p-type carriers and a negative value indicates n-type carriers.
| ( cm3) | ( cm3) | (cm2V-1s-1) | (cm2V-1s-1) | |
| Fitted values | ||||
| Upper bounds | ||||
| Lower bounds |
Figure S2 illustrates the two-band fitting of the high-field and curves measured at base temperature. To estimate the uncertainty range of the fitting parameters, we vary around its optimal value and then optimize the other parameters to minimize the residuals (defined as ) between the measured and fitted curves. We then compare the standard deviation of the residuals with its minimized value obtained under the best-fit condition. The boundary of the parameter range is defined as the point at which the standard deviation of the residuals reaches ten times its best-fit value. The fitted values and uncertainty ranges of all fitting parameters are listed in Table S1.
Appendix S4 Procedure for Classical Magnetoresistance Background Subtraction
As described in the main text, the raw data contain a large positive classical magnetoresistance (MR) in addition to a smaller quantum correction. To isolate the quantum contribution, we fitted the two-band Drude model to the high-field part of the data, thereby obtaining the corresponding classical background . The preliminary background-subtracted curve is then defined as
Since this subtraction procedure uses the high-field region as the reference point, does not naturally vanish at . To make the result directly comparable with theoretical descriptions of quantum corrections (such as the Kondo effect), which characterize the field-induced relative change of resistivity from its zero-field value, we further shift the curve by subtracting :
This procedure ensures that , consistent with the theoretical framework. The resulting curves shown in Figs. 3(e)–(f) of the main text therefore represent the relative change of resistivity under magnetic field after removing the classical background. For illustration, Fig. S3 shows an example of this background subtraction process using the data from Device A.